OpenStudy (anonymous):

help me in finding the derivative of the function below pls y = 7x^3(x^3-3)^5 / (4x+1)^3(7-x^5)^6

OpenStudy (anonymous):

as there is a division try applying the quotient rule (vdu-udv/v^2) the numbers are very awkward youll probably want to use a calculator, or derive it with that if your calculator can

OpenStudy (anonymous):

entering it how you posted it into mine i got -(60x^16+12x^15-1080x^13-225x^12+7560x^10+1620x^9-25920x^7-5670x^6+43740x^4+9720x^3-29160x-6561)/(16807x^28(4x+1)^4)

OpenStudy (anonymous):

@Taplin44 i attached a picture of the function so you can understand clearly. pls help me

OpenStudy (anonymous):

The question asks for dy/dx yes ? if so the verrrry long string of numbers i posted is the correct answer note that 16807x^28(4x+1)^4 is all the denominator to the rest

OpenStudy (anonymous):

@Taplin44 how bout this picture? is it clear for you?

OpenStudy (amistre64):

just pull the bottom up and run a 4 term product rule ... abcd \(\color{red}{\to}a'bcd+ab'cd+abc'd+abcd'\)

OpenStudy (anonymous):

@amistre64 do i have to use quotient rule, product rule, and chain rule?

OpenStudy (amistre64):

id just use the product rule on something this messy

OpenStudy (amistre64):

\[y=7x^3~(x^3-3)^5~(4x+1)^{-3}~(7-x^5)^{-6}\]

OpenStudy (anonymous):

The picture is clear. It is just alot of arithmetic to do by hand hence why i evaluated with a calculator for you The quotient or product rule will work. Using the product rule requires you to bring up the denominator but as @amistre64 said, is going to be a lot easier

OpenStudy (amistre64):

the results will still be a long string tho :) and you would be able to simplify it by factoring out what amounts to the square of the bottom

OpenStudy (anonymous):

@taplin44 well our teacher requires us to use the quotient rule, product rule, and chain rule

OpenStudy (amistre64):

the quotient rule is pretty well obsolete since dividing is just another way of multiplying. The rule is a throw back to the limit method i believe

OpenStudy (amistre64):

you are still applying the product rule, the power rule, and the chain rule in this setup

OpenStudy (anonymous):

okaay thank u :)

OpenStudy (anonymous):

@Taplin44 @amistre64 how to find the derivative of the numerator? help

OpenStudy (amistre64):

the numerator is just a product ... ab \(\to\) a'b + ab'

OpenStudy (amistre64):

let a = 7x^3 let b=(x^3-3)^5 define their own prospective derivatives and fill in the rule ....

OpenStudy (anonymous):

@amistre64 @Taplin44 how to simplify the attached? help

OpenStudy (anonymous):

@rose21 here!

OpenStudy (rose21):

yea I see lol

OpenStudy (anonymous):

@rose21 attached is the equation. find the derivative. help me

OpenStudy (anonymous):

you have to use logarithemc diffrentation

OpenStudy (anonymous):

@Ahmad1 we haven't encountered that topic yet. all i know is to find the derivative by using quotient rule, product rule, and chain rule

OpenStudy (anonymous):

trust me I'm telling you the idea for such questions , it's simple by the way \[\ln(y)=3\ln(7x)+5\ln(x^3-3)-3\ln(4x+1)-6\ln(7-x^5)\], now differntiate both sides , the LHD is dy/dx *1/y , so multiply by y to get y prime , i.e\[dy/dx=y*(RHS) \prime\] which is easy to find

OpenStudy (anonymous):

im sorry but idk what that is

OpenStudy (anonymous):

don't you know the logarithmic function ? I have just used its properties

OpenStudy (anonymous):

no sorry. we haven't encountered that yet.

OpenStudy (anonymous):

okay no problem ..

OpenStudy (anonymous):

@UsArmy3947 here!

OpenStudy (anonymous):

@UsArmy3947 find the derivative

OpenStudy (anonymous):

@Ahmad1 how to simplify the attached?

OpenStudy (anonymous):

@rose21 how to simplify the attached?

OpenStudy (usarmy3947):

ok simplify the attachment on top?

OpenStudy (anonymous):

@UsArmy3947 this one

OpenStudy (usarmy3947):

ok thx

OpenStudy (usarmy3947):

ok i cant help u but ill ask my friend ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is this the derivative of the function?

OpenStudy (usarmy3947):

@thomaster @shkrina @robtobey @elisuzsmith @razor99 @terenzreignz @tshark14 @Tron_Cat @YourMentor

OpenStudy (anonymous):

@Ashleyisakitty @andriod09 @allie_bear22 @ankit042 @Ashja @ashley_f97

OpenStudy (usarmy3947):

im sorry if i couldn't help u

OpenStudy (anonymous):

it's okay :)

OpenStudy (amistre64):

its prolly best not to try to "simplify" it; and just to keep it in its "factored" form. at least thats my opinion

OpenStudy (anonymous):

@amistre64 u mean that's the final answer?

OpenStudy (usarmy3947):

@amistre64 @Hero @dumbcow @dan815 @dmezzullo @DebbieG @dlapointe25 @Deivyneee @andriod09 @allie_bear22 @ankit042 @Ashleyisakitty @AnElephant @bashirk @iceicebaby

OpenStudy (amistre64):

\[y=7x^3~(x^3-3)^5~(4x+1)^{-3}~(7-x^5)^{-6}\] \[y'=\\ ~~~~7(3)x^2~(x^3-3)^5~(4x+1)^{-3}~(7-x^5)^{-6}\\ +35(3)x^5(x^3-3)^4~(4x+1)^{-3}~(7-x^5)^{-6}\\ -21(4)x^3~(x^3-3)^5~(4x+1)^{-4}~(7-x^5)^{-6}\\ +42(5)x^7~(x^3-3)^5~(4x+1)^{-3}~(7-x^5)^{-7}\] then factoring out a : \((4x+1)^{-6}~(7-x^5)^{-12}\) \[y'=[(4x+1)^{-6}~(7-x^5)^{-12}]\\ [~~~~7(3)x^2~(x^3-3)^5~(4x+1)^{3}~(7-x^5)^{6}\\ +35(3)x^5(x^3-3)^4~(4x+1)^{3}~(7-x^5)^{6}\\ -21(4)x^3~(x^3-3)^5~(4x+1)^{2}~(7-x^5)^{6}\\ +42(5)x^7~(x^3-3)^5~(4x+1)^{3}~(7-x^5)^{4}]\]

OpenStudy (amistre64):

i spose the squaring of the bottom is a bit much ... we could simply factor out the higest negative degrees instead for a more simplified form

OpenStudy (amistre64):

\[y'=[(4x+1)^{-4}~(7-x^5)^{-7}]\\ [~~~~7(3)x^2~(x^3-3)^5~(4x+1)~(7-x^5)\\ +35(3)x^5(x^3-3)^4~(4x+1)~(7-x^5)\\ -21(4)x^3~(x^3-3)^5~~(7-x^5)\\ +42(5)x^7~(x^3-3)^5~(4x+1)]\]

OpenStudy (amistre64):

factoring out the smallest exponents of the top gets us: \[y'=[x^2(x^3-3)^4]\\ [~~~~7(3)~(x^3-3)~(4x+1)~(7-x^5)\\ +35(3)x^3~(4x+1)~(7-x^5)\\ -21(4)x~(x^3-3)~~(7-x^5)\\ +42(5)x^5~(x^3-3)~(4x+1)]\\ -----------------\\ ~~~~~~~~~~(4x+1)^{4}~(7-x^5)^{7}\] and theres some constants we could pull as well